OpenStudy (anonymous):

A firefighter directs a horizontal stream of water toward a fire at an angle of 40° measured from the horizontal. The fire hydrant and hose deliver 93 gallons per minute (1 gal/min = 6.31 cross product 10−5 m3/s), and the nozzle projects the water with a speed of 6.7 m/s. How much horizontal force must the firefighter exert on the hose to keep it stationary?

OpenStudy (irishboy123):

F = dp/dt [Newton's original formulation] = d(mv)/dt = v•dm/dt + m•dv/dt dV/dt is just a proxy for dm/dt. introduce a placeholder for density. you will need to call it µ or the equations just won't work. µ dV/dt = dm/dt. thus, dV/dt = (1/µ ) * dm/dt. these need to be resolved in the x and y directions.